Phenol(C6H5OH) is commonly used to make a buffer with Sodium Phenolate (C6H5ONa) . Given that the Ka = 1.29*10^-10, and a buffer was created with 0.25g of Phenol and 0.50g Sodium Phenolate. Determine the pH after the addition of 2mL of 0.10M HCI.


1
Expert's answer
2021-04-27T07:48:31-0400

Molar Mass of Phenol = 94.11

0.25/94.11 = 0.00266M

Molar Mass of sodium phenolate = 116.09

= 0.5/116.09

= 0.00431M

Molar Mass of HCl 36.458

36.458 × 0.1

= 3.6458 ×2

= 7.2916/36.458

= 0.2moles

0.2/0.75 = X/1L

= 0.15 of H+

PH =14+ -log(0.15)

= 14+0.82

= 14.82